# Structural and Dynamical Crossovers in Dense Electrolytes

### v1

@@ Abstract
Electrostatic interactions fundamentally govern the structure and transport of electrolytes. In concentrated electrolytes, however, electrostatic and steric correlations, together with ion–solvent coupling, give rise to complex behavior, such as underscreening, that remains challenging to explain despite extensive theoretical effort. Using molecular dynamics simulations of primitive electrolytes with and without space-filling solvent particles, we elucidate the structural and dynamical crossovers and their connection that emerge with increasing salt concentration. Explicit-solvent electrolytes exhibit a screening transition from a charge-dominated dilute regime to a density-dominated concentrated regime, accompanied by dynamical crossovers in ion self-diffusion and ion-pair lifetimes. These dynamical crossovers display a marked discontinuity, unlike the smoother variation of the screening crossover, which originates from short-range ion–counterion structures. Despite the pronounced growth of ionic clusters, their percolation transition does not appear to be directly coupled to the onset of these crossovers. Both structural and dynamical behaviors are found to depend sensitively on ion–solvent coupling: implicit-solvent electrolytes exhibit a screening transition between two charge-dominated regimes, accompanied by qualitatively distinct dynamical behavior. Finally, we demonstrate that the diffusion-corrected ion-pair lifetime provides a consistent descriptor linking ionic structure and dynamics across electrolyte systems.
@@ I Introduction
[…]
@@
We identify dynamic crossovers in ion self-diffusion and ion-pair dynamics that accompany the screening crossover and can be unified by scaling with the ratio of the mean ion size to the Debye screening length, as used in the SFB measurements. The nature of the screening crossover depends sensitively on explicit-solvent packing: explicit-solvent systems transition from a charge-dominated dilute regime to a density-dominated concentrated regime, whereas implicit-solvent systems undergo a transition between two charge-dominated regimes. Although enhanced ion–ion correlations promote a percolation transition in ionic clustering, we find that this percolation is not directly coupled to either screening or dynamical crossovers; instead, it correlates more strongly with ion density. The dynamical crossovers exhibit a marked discontinuity, linked to qualitative changes in short-range ion–counterion structure. Finally, we identify the diffusion-corrected ion-pair lifetime as a unified microscopic descriptor that consistently bridges ionic structure and transport across electrolyte models.
@@ II.2 Computation of observables
The normalization constant $N_{s}$ ensures $\sum_{s}P(s)=1$. This analysis is analogous to the scaling approaches used in critical phenomena [82, 83, 59]. Two characteristic length scales, $\xi_{s}$ and $\gamma_{s}$, describe different parts of the distribution: the former governs the exponential decay, while the latter characterizes the power-law tail. Their relative contributions change across the percolation (or gelation) transition. In the pre-gel regime, the exponential part dominates, corresponding to small, localized ion clusters. At the gelation point, $P(s)$ reduces to a pure power-law distribution, implying the emergence of a scale-free ionic network spanning the simulation box. In the post-gel regime, the percolating ionic network coexists with small clusters.
@@
In this work, we focused on the self-diffusion ($\alpha=\beta$), where the MSD of each species grows linearly with time within the simulation window (see Sec. S7 in the Supplementary Material (SM) [75]). Thanks to the charge inversion symmetry, we used $D_{\mathrm{ion}}=(D_{++}+D_{--})/2=D_{++}=D_{--}$. Using $D_{\alpha\beta}$, we also estimated the characteristic diffusion timescale over which an ion diffuses, on average, a distance comparable to its diameter: $\tau_{\mathrm{diff}}=\frac{\sigma^{2}}{6D_{\mathrm{ion}}}$ [88].
@@ III Results and Discussion
Figure 2(C) summarizes our results for both LJ and WCA electrolytes, which reproduce the electrostatic screening crossovers observed in SFB experiments, yet yield underestimated scaling exponents ($\alpha\approx 1.5$22). Consistent with previous bulk molecular simulations [33, 34, 29] and classical DFT studies [35, 36, 30], this underestimation leaves the physical basis of the experimental scaling unresolved, a discrepancy that persists across both chemically detailed models and coarse-grained systems such as primitive electrolytes. Recent studies [67, 93, 52] have further suggested that confinement- or interface-induced effects, including enhanced ion association, may contribute to the anomalous cubic scaling observed in SFB experiments, but investigating these effects is beyond the scope of the present study.
@@ III.2 Formation of a percolating ionic network is not directly coupled to the structural crossover
Figure 3: Percolation transition of ionic clusters in explicit-solvent electrolytes. (A) Evolution of ionic clusters with salt concentration $c_{\mathrm{salt}}$. The composition of each cluster is defined by the number pair ($_{+}, nn_{-}$). (B) Size distribution $P(s)$ (Eq. 9) of ionic clusters with $s=n_{+}+n_{-}$. $c_{\mathrm{salt}}$ increases with the color changing from purple to green to red. The black dotted line indicates the power-law decay at the percolation threshold, $c_{\mathrm{gel}}\approx 0.063~\sigma^{-3}$, $P(s)\sim s^{-1.85}$. Panels (A) and (B) show results for explicit-solvent LJ electrolytes with $\varepsilon_{s}=0.2$. (C) Dependence of the exponential decay length $\xi_{s}$, extracted from $P(s)$, on salt concentration $c_{\mathrm{salt}}$ across all explicit-solvent electrolytes. In panel (C), the colors and markers are consistent with those in Fig. 1.
In an attempt to examine a connection between the screening crossover and the formation of a percolating ionic network, we quantify the growing size of ionic clusters with increasing $c_{\mathrm{salt}}$ due to increased ion-ion correlations (Fig. 3). In defining an ionic cluster, a pair of a cation and an anion belongs to the same cluster if their interionic distance is within the cutoff distance $r_{cut}=1.5~\sigma$ taken from the first minimum of $g_{+-}(r)$ (see Fig. S14, S15 in the SM [75]). Due to the charge symmetry in the systems of interest, the numbers of constituent cations $n_{+}$ and anions $n_{-}$ of an ionic cluster are the same. We characterize the ionic clusters by their size distribution $P(s)$ (Fig. 3 and Tab. 1). Two characteristic length scales ($\gamma_{s}$ and $\xi_{s}$ in Eq. 9) describe different parts of the distribution: the former governs the exponential decay, while the latter characterizes the power-law tail. At sufficiently low $c_{\mathrm{salt}}$, $P(s)$ is well described by the exponential part alone with its decay length $\xi_{s}$. Far below the threshold salt concentration (the pre-gel regime), $\gamma_{s}$ increases with $c_{\text{salt}}$ as more ionic clusters form, while the contribution of the power-law tail grows with $\gamma_{s}$, approaching the threshold. At the threshold concentration $c_{\mathrm{gel}}$, $P(s)$ follows a pure power-law distribution, indicating the emergence of a scale-free ionic network spanning the simulation box. This threshold corresponds to the gelation or percolation point ($c_{\text{salt}}\approx 0.063~\sigma^{-3}$ for LJ electrolytes with $\varepsilon_{s}=0.2$). In the post-gel regime, the power-law tail diminishes with increasing salt concentration, as a larger fraction of ions participates in the spanning network, thereby reducing the relative contribution of small ionic clusters.
@@
Despite the intriguing $c_{\text{salt}}$ dependence of $P(s)$, the formation of an ionic network does not appear to be strongly correlated with the screening crossover (Fig. 3(B)). Instead, the emergence of a percolating network seems to be primarily governed by the spatial proximity of ions, as it occurs even in relatively dilute electrolytes (e.g., $\varepsilon_{s}=5$), with gelation taking place at a similar transition concentration, $c_{\mathrm{gel}}$ (Tab. 1). Therefore, the screening crossover appears to be decoupled from the ionic network formation in these solvent-dense concentrated electrolytes. We note that the emergence of a percolating network, in general, provides a structural basis for strong correlations in ionic motion [59, 61, 62, 63]. Recent studies [64, 65] have further suggested its relation to the cubic scaling observed in SFB experiments. In connection with the growing ionic domains, two competing length scales were introduced—the screening length, defined in terms of renormalized charges, and the average cluster size, which increases up to the percolation threshold—whose larger value is proposed to govern the scaling behavior.
@@ IV Conclusions
Using molecular dynamics simulations of model electrolytes with and without explicit solvent, we systematically investigated the relationships between structural and dynamical correlations in dense ionic fluids across a wide range of salt concentrations. We find that the electrostatic screening crossover depends sensitively on ion–solvent coupling. In explicit-solvent electrolytes, the crossover occurs between a charge-dominated dilute regime and a density-dominated concentrated regime, whereas in implicit-solvent electrolytes, it occurs between two charge-dominated regimes—from dilute to concentrated. Both explicit- and implicit-solvent systems display dynamical crossovers in ion self-diffusion and ion-pair lifetimes that are coupled to the screening transition, although the detailed dynamical behaviors differ depending on the presence of solvent. The notable discontinuity across the dynamical transition is well encoded in the ion–counterion short-range structures. Ionic clusters grow with increasing salt concentration and eventually form a percolating network beyond a threshold concentration. However, the percolation transition is not strongly coupled to the onset of either the structural or dynamical crossovers and appears instead to be governed primarily by salt concentration. This decoupling suggests that spatial proximity alone is insufficient to explain the connection between ionic clustering and transport in these model systems.
@@
Supplementary material
Supplementary Material provides detailed summaries and fitting results of structural, dynamical, and transport properties across all electrolyte conditions examined in this study. These include charge–charge, density–density, solvent–solvent, cation–anion, cation-solvent correlation functions, cluster-size distributions, mean-squared displacements of ions and solvent, ion-pair survival functions, and ionic conductivity.
@@ Acknowledgment
This work was supported by a New Faculty Research Grant of Pusan National University, 2025. This research was supported by the Regional Innovation System & Education(RISE) program through the Jeju RISE center, funded by the Ministry of Education(MOE) and the Jeju Special Self-Governing Province, Republic of Korea.(2025-RISE-17-001)
@@
The data that support the findings of this article are openly available upon acceptance of the manuscript [101].

### v2

@@ Abstract
Electrostatic interactions fundamentally govern the structure and transport of electrolytes. In concentrated electrolytes, however, electrostatic and steric correlations, together with ion–solvent coupling, give rise to complex behavior, such as underscreening, that remains challenging to explain despite extensive theoretical effort. Using molecular dynamics simulations of primitive electrolytes with and without space-filling solvent particles, we elucidate the structural and dynamical crossovers and their connection that emerge with increasing salt concentration. Explicit-solvent electrolytes exhibit a screening transition from a charge-dominated dilute regime to a density-dominated concentrated regime, accompanied by dynamical crossovers in ion self-diffusion and ion-pair lifetimes. These dynamical crossovers display a marked discontinuity, unlike the smoother variation of the screening crossover, which originates from short-range ion–counterion structures. The pronounced growth of ionic clusters leads to a percolation transition only at higher concentrations, whereas the onset of the structural and dynamical crossovers is associated primarily with local ion pairing or small aggregates. Both structural and dynamical behaviors are found to depend sensitively on ion–solvent coupling: implicit-solvent electrolytes exhibit a screening transition between two charge-dominated regimes, accompanied by qualitatively distinct dynamical behavior. Finally, we demonstrate that the diffusion-corrected ion-pair lifetime provides a consistent descriptor linking ionic structure and dynamics across electrolyte systems.
@@ I Introduction
In addition to the structural features, understanding and predicting the dynamic behavior of concentrated electrolytes remain equally challenging. Theoretical approaches, such as mean-spherical approximation (MSA) [37, 38] and stochastic density functional theory (sDFT) [39, 40, 41], have extended the classical Debye-Hückel-Onsager theory [42, 43] by incorporating steric ion-ion interactions. Recent applications of sDFT [44, 41, 45] have reported encouraging agreement with experimental trends in the transport properties of moderately concentrated aqueous electrolytes. Nevertheless, a subsequent study [46] indicated that the quantitative accuracy of these predictions may be limited by compensating effects among different theoretical contributions. However, a unified microscopic understanding that consistently connects ion dynamics and electrostatic screening remains elusive, seeking to establish a direct connection between structure and dynamics. That is, how microscopic ion and solvent organization governs macroscopic transport properties, such as electrical conductivity and diffusion.
@@
We identify dynamic crossovers in ion self-diffusion and ion-pair dynamics that accompany the screening crossover and can be unified by scaling with the ratio of the mean ion size to the Debye screening length, as used in the SFB measurements. The nature of the screening crossover depends sensitively on explicit-solvent packing: explicit-solvent systems transition from a charge-dominated dilute regime to a density-dominated concentrated regime, whereas implicit-solvent systems undergo a transition between two charge-dominated regimes. Although enhanced ion–ion correlations promote a percolation transition in ionic clustering, we find that this percolation is not directly coupled to either screening or dynamical crossovers. Instead, it correlates more strongly with ion density, reflecting that percolation requires sufficiently extensive aggregation to form a system-spanning network and therefore typically occurs only at higher salt concentrations, whereas the onset of underscreening is associated with the emergence of local ion pairing or small aggregates. The dynamical crossovers exhibit a marked discontinuity, linked to qualitative changes in short-range ion–counterion structure. Finally, we identify the diffusion-corrected ion-pair lifetime as a unified microscopic descriptor that consistently bridges ionic structure and transport across electrolyte models.
@@ II.2 Computation of observables
The normalization constant $N_{s}$ ensures $\sum_{s}P(s)=1$. This analysis is analogous to the scaling approaches used in critical phenomena [83, 84, 60]. Two characteristic length scales, $\xi_{s}$ and $\gamma_{s}$, describe different parts of the distribution: the former governs the exponential decay, while the latter characterizes the power-law tail. Their relative contributions change across the percolation (or gelation) transition. In the pre-gel regime, the exponential part dominates, corresponding to small, localized ion clusters. At the gelation point, $P(s)$ reduces to a pure power-law distribution, implying the emergence of a scale-free ionic network spanning the simulation box. In the post-gel regime, the percolating ionic network coexists with small clusters. Additional analysis of ionic clusters, including association patterns and comparison to the Cayley-tree-based ionic gelation theory [60, 85, 63, 86, 31], is provided in the SM [76].
@@
In this work, we focused on the self-diffusion ($\alpha=\beta$), where the MSD of each species grows linearly with time within the simulation window (see Sec. S7 in the Supplemental Material (SM) [76]). Thanks to the charge inversion symmetry, we used $D_{\mathrm{ion}}=(D_{++}+D_{--})/2=D_{++}=D_{--}$. Using $D_{\alpha\beta}$, we also estimated the characteristic diffusion timescale over which an ion diffuses, on average, a distance comparable to its diameter: $\tau_{\mathrm{diff}}=\frac{\sigma^{2}}{6D_{\mathrm{ion}}}$ [91].
@@ III Results and Discussion
Figure 2(C) summarizes our results for both LJ and WCA electrolytes, which reproduce the electrostatic screening crossovers observed in SFB experiments, yet yield underestimated scaling exponents ($\alpha\approx 1.5$22). Notably, the range of measured $\lambda_{Z}$ extends only up to $\sim 10~\lambda_{D}$, whereas anomalous underscreening in SFB experiments has been reported to reach $\sim 100~\lambda_{D}$. Consistent with previous bulk molecular simulations [33, 34, 29] and classical DFT studies [35, 36, 30], this underestimation leaves the physical basis of the experimental scaling unresolved, a discrepancy that persists across both chemically detailed models and coarse-grained systems such as primitive electrolytes. Recent studies [68, 96, 53] have further suggested that confinement- or interface-induced effects, including enhanced ion association, may contribute to the anomalous cubic scaling observed in SFB experiments, but investigating these effects is beyond the scope of the present study.
@@ III.2 Formation of a percolating ionic network is not directly coupled to the structural crossover
Figure 3: Percolation transition of ionic clusters in explicit-solvent electrolytes. (A) Evolution of ionic clusters with salt concentration $c_{\mathrm{salt}}$. The composition of each cluster is defined by the number pair ($_{+}, nn_{-}$). (B) Size distribution $P(s)$ (Eq. 9) of ionic clusters with $s=n_{+}+n_{-}$. $c_{\mathrm{salt}}$ increases with the color changing from purple to green to red. The black dotted line indicates the power-law decay at the percolation threshold, $c_{\mathrm{gel}}\approx 0.063~\sigma^{-3}$, $P(s)\sim s^{-1.85}$. Panels (A) and (B) show results for explicit-solvent LJ electrolytes with $\varepsilon_{s}=0.2$. (C) Dependence of the exponential decay length $\xi_{s}$, extracted from $P(s)$, on salt concentration $c_{\mathrm{salt}}$ across all explicit-solvent electrolytes. In all subpanels of panel (A), the probability at the point $(n_{+},n_{-})=(0,0)$ is zero; the apparent non-zero value indicated by the dark blue color arises solely from the finite binning used in constructing the heat map. In panel (C), the colors and markers are consistent with those in Fig. 1.
In an attempt to examine a connection between the screening crossover and the formation of a percolating ionic network, we quantify the growing size of ionic clusters with increasing $c_{\mathrm{salt}}$ due to increased ion-ion correlations (Figs. 3). In defining an ionic cluster, a pair of a cation and an anion belongs to the same cluster if their interionic distance is within the cutoff distance $r_{cut}=1.5~\sigma$ taken from the first minimum of $g_{+-}(r)$ (see Figs. S10 and S14-S15 in the SM [76]). Due to the charge symmetry in the systems of interest, the numbers of constituent cations $n_{+}$ and anions $n_{-}$ of an ionic cluster are the same. We characterize the ionic clusters by their size distribution $P(s)$ (Fig. 3 and Tab. 1). Two characteristic length scales ($\gamma_{s}$ and $\xi_{s}$ in Eq. 9) describe different parts of the distribution: the former governs the exponential decay, while the latter characterizes the power-law tail. At sufficiently low $c_{\mathrm{salt}}$, $P(s)$ is well described by the exponential part alone with its decay length $\xi_{s}$. Far below the threshold salt concentration (the pre-gel regime), $\gamma_{s}$ increases with $c_{\text{salt}}$ as more ionic clusters form, while the contribution of the power-law tail grows with $\gamma_{s}$, approaching the threshold. At the threshold concentration $c_{\mathrm{gel}}$, $P(s)$ follows a pure power-law distribution, indicating the emergence of a scale-free ionic network spanning the simulation box. This threshold corresponds to the gelation or percolation point ($c_{\text{salt}}\approx 0.063~\sigma^{-3}$ for LJ electrolytes with $\varepsilon_{s}=0.2$). In the post-gel regime, the power-law tail diminishes with increasing salt concentration, as a larger fraction of ions participates in the spanning network, thereby reducing the relative contribution of small ionic clusters.
@@
Despite the intriguing $c_{\text{salt}}$ dependence of $P(s)$, the formation of an ionic network does not appear to be strongly correlated with the screening crossover (Fig. 3(B)). Instead, the emergence of a percolating network seems to be primarily governed by the spatial proximity of ions, as it occurs even in relatively dilute electrolytes (e.g., $\varepsilon_{s}=5$), with gelation taking place at a similar transition concentration, $c_{\mathrm{gel}}$ (Tab. 1). Therefore, the screening crossover appears to be decoupled from the ionic network formation in these solvent-dense concentrated electrolytes. This decoupling can be rationalized by the distinct physical requirements of the two phenomena: percolation requires sufficiently extensive aggregation to form a system-spanning network and thus typically occurs only at higher salt concentrations, whereas the onset of underscreening is associated with the emergence of local ion pairing or small aggregates. We note that the emergence of a percolating network, in general, provides a structural basis for strong correlations in ionic motion [60, 62, 63, 64]. Recent studies [65, 66] have further suggested its relation to the cubic scaling observed in SFB experiments. In connection with the growing ionic domains, two competing length scales were introduced—the screening length, defined in terms of renormalized charges, and the average cluster size, which increases up to the percolation threshold—whose larger value is proposed to govern the scaling behavior.
The Cayley-tree-based ionic gelation theory [60, 85, 63, 86, 31] provides a theoretical framework for describing ionic clusters in concentrated electrolytes at the mean-field level, thereby linking structural changes—particularly gelation—to other physical properties. This framework has been successfully applied to a range of concentrated electrolytes, including (diluted) ionic liquids and their electrochemical interfaces. A key assumption of the theory is that ions form Cayley-tree clusters among other possible morphologies, such as chains, rings, and branched ionic aggregates. In contrast, owing to the simplicity of our electrolyte models, which employ isotropic interaction potentials between all ions and solvent particles, the resulting association patterns are more complex than ideal Cayley-tree-like structures (see Figs. S30-S31 in the SM [76]), particularly owing to the presence of co-ion associations and intra-cluster loops. These deviations, together with ambiguities in defining the ion-association functionality - a crucial input parameter of the theory - hinder the theory from accurately locating the gelation threshold $c_{\text{gel}}$ of our simple models over the estimated counterion-association functionality range of 2-5.
Similar limitations of the ionic gelation theory have been reported previously; for example, theoretical predictions of gelation points were found to be systematically lower than those estimated from simulations for nonaqueous LiPF6 and LiBF4 electrolytes, a discrepancy likewise attributed to deviations from ideal Cayley-tree clusters [63]. Nevertheless, we find that below $c_{\text{gel}}$ the ionic gelation theory captures the counterion association constant reasonably well, yielding values of comparable order of magnitude to those observed in simulations when an appropriate ion-association functionality is chosen (see Fig. S32 in the SM [76]). A detailed analysis and discussion of ionic association, together with its comparison with the ionic gelation theory, are provided in Sec. S10 of the SM [76].
@@ IV Conclusions
Using molecular dynamics simulations of model electrolytes with and without explicit solvent, we systematically investigated the relationships between structural and dynamical correlations in dense ionic fluids across a wide range of salt concentrations. We find that the electrostatic screening crossover depends sensitively on ion–solvent coupling. In explicit-solvent electrolytes, the crossover occurs between a charge-dominated dilute regime and a density-dominated concentrated regime, whereas in implicit-solvent electrolytes, it occurs between two charge-dominated regimes—from dilute to concentrated. Both explicit- and implicit-solvent systems display dynamical crossovers in ion self-diffusion and ion-pair lifetimes that are coupled to the screening transition, although the detailed dynamical behaviors differ depending on the presence of solvent. The notable discontinuity across the dynamical transition is well encoded in the ion–counterion short-range structures. Ionic clusters grow with increasing salt concentration and eventually form a percolating network beyond a threshold concentration. However, the percolation transition is not strongly coupled to the onset of either the structural or dynamical crossovers and appears instead to be governed primarily by salt concentration. This decoupling can be rationalized by the distinct physical requirements of the two phenomena: percolation requires sufficiently extensive aggregation to form a system-spanning network and thus typically occurs only at higher salt concentrations, whereas the onset of underscreening reflects the emergence of local ion pairing or small aggregates. These observations also suggest that spatial proximity alone is insufficient to explain the connection between ionic clustering and transport in these model systems.
@@
Supplemental material
Supplemental Material provides detailed summaries and fitting results of structural, dynamical, and transport properties across all electrolyte conditions examined in this study. These include charge–charge, density–density, solvent–solvent, cation–anion, cation-solvent correlation functions, cluster-size distributions, mean-squared displacements of ions and solvent, ion-pair survival functions, ionic conductivity, and a detailed analysis of ionic associations.
@@ Acknowledgment
This work was supported by the National Supercomputing Center with supercomputing resources including technical support(KSC-2025-CRE-0142). This research was supported by the Regional Innovation System & Education(RISE) program through the Jeju RISE center, funded by the Ministry of Education(MOE) and the Jeju Special Self-Governing Province, Republic of Korea(2025-RISE-17-001).
@@
The data that support the findings of this article are openly available [105].